### 14 Dec 2023 v3

@@ 4.4 Exploratory data analysis (EDA) & data reduction (DR)
In this process, the target or dependent column and independent features are obtained. The DR, EDA, and clustering techniques reduce runtime and space during the deep-learning modeling phase. DR can be employed to decrease the size of a dataset while still keeping the data’s original integrity. In our framework, we suggest performing Feature Selection and Feature Extraction simultaneously after selecting the target column and finding redundant features; then, Discretisation may be performed if necessary. Then the dataset will be ready for further analysis and model training.
@@ 4.4.1 Identifying redundant features
Covariance and correlation: In statistics, covariance refers to the amount that two features or factors change in tandem whose value lies in the (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) range. Positive covariance indicates they move in the same direction. Negative covariance means that any features are greater than the mean, and others are less than vice-versa. Zero covariance means features may be independent under a certain hypothesis. On the other hand, correlation analysis is a widely used dimensionless measurement ranging from $-1$ to $+1$ to discover redundancies in numerical features that evaluate and quantify the relationship intensity. The features are positively correlated for correlation values greater than zero (0); for zero, they are independent; for less than zero, they are negatively correlated [46]. Covariance and correlation are directly proportional to each other. In numeric feature selection, correlation is better to use, as correlation analysis is scaled [-1, 1], but the covariance range is indefinite (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). We should choose correlation for better interpretation. Changes in location, size, or scale have no effect on correlation. However, both of them are limited to only being able to identify linearity.
$\chi^{2}$ correlation: The $\chi^{2}$ (Chi-Square) test is often used when dealing with nominal features and finite value-set. We can use the $\chi^{2}$ test to see whether there is any link between the values of two nominal features, where a probability table with joint events is established. If $\alpha$ (significance level) is less than the estimated one (or the $\chi^{2}$-value (calculated) ¿ table value), the null hypothesis gets discarded, and the two features can be said to be correlated statistically [46]. SC Analysts must remember that the $\chi^{2}$-test does not tell much about the strength of the relationship between two features. The $\chi^{2}$-test offers advantages such as resilience regarding data distribution, computational simplicity, extensive information produced from the test, utilization for investigations where parametric criteria cannot be satisfied, and scalability in processing data from two and multiple-group research. The drawbacks are sample size constraints and difficulty in comprehension when there are many (¿20) features.
@@ 4.4.2 Feature selection(FS)
  • In certain circumstances, FS results will still include significant important features which may obstruct using complicated training strategies.
Leading methods: In order to create FS techniques by combining a feature evaluation score and a cutting criterion, [3] recommended that functions based on information principle produce better accuracy, not suggesting any universal cutting condition. However, those independent of the metric perform best, and outcomes differ across models, and for each kind of model, wrapper techniques were recommended to avoid this effect.
@@
We can use supervised FS if the data has class labels; else, unsupervised FS is the best option. This approach generally maximizes clustering efficiency or the FS based on correlation, feature dependency, and priority. The primary premise is to eliminate features that bring almost no value beyond what is already provided by the existing features in the system. [94] suggested using feature dependency/similarity to reduce redundancy without needing a search procedure. An information compression metric called the maximum information compression index governs the clustering partitioning process, which uses features as the measure of similarity. Forward orthogonal search (FOS) is another unsupervised FS approach that aims to maximize the total reliance on the data to find relevant features [141]. Without compromising performance in clustering, the unsupervised FS used the Random Cluster Ensemble framework to compress the set of features by roughly 1/100 of its initial dimensions [32]. When compared to well-known classifications, precision/recall analyses revealed that feature weighting was highly successful in discovering the most suitable clusters [96].
@@ 4.4.3 Feature extraction
Principal component analysis (PCA): It targets to analyze a collection of features’ variance-covariance patterns employing a few linear combinations and seeks the optimal $k$ number of $n-dimensional$ orthogonal vectors for data description, where $k\geq n$. Accounting for the most critical percentage of the discrepancy in the original dataset, the principal component (the first derived feature) is produced in decreasing order of contribution. Typically, for containing $\geq$ 95% variance, just the top few principal components are retained. PCA is beneficial when many independent variables correlate with one another [70]. Principal-component is quick and comprehensive and ensures a solution is found for all datasets [44].
@@ 4.5.2 Model fit and train
Two factors determine the proportions of these three sets: the number of data samples and training models. Some models require significant training data; therefore, the model should be tuned for more extensive training sets in this scenario. Models with fewer hyperparameters will be easier to validate and tune, allowing a small validation set size. An extensive validation set will be beneficial if the model contains considerably more significant hyperparameters. There will be no requirement for a validation set if the model has no hyper-parameters or is challenging to adjust.
When using k-fold CV, the train-test dataset splitting is repeated for k-times, with each new set being given a shot at becoming the hold-out set. Time-series data cannot be used with k-fold CV directly since they believe that there lies no connection between the rows and that they are all separate instances. For time-series data, instances’ time horizon prevents arbitrarily dividing them into clusters. Instead, data should be segmented, and the chronological sequence of instances maintained. The term backtesting is used in time-series forecasting to describe the technique of evaluating models using past data. In meteorology, this is regarded as ’hindcasting’ rather than ’forecasting.’
@@ 4.5.3 Hyperparameter tuning
  • It enhances the opportunity of reproducing the ML process
@@ 4.5.5 Top forecasting models
Adaptive Neuro–Fuzzy Inference System (ANFIS) [55]
@@
For the demand forecasting procedure, [149] evaluated statistical models, RBFNN (Radial Basis Function NNs), and winter models with SVM. According to their conclusion, the efficiency of SVM outperforms other algorithms by about a mean MAPE outcomes threshold of 7.7%. [30] demonstrated a new AI-utilized forecasting approach evaluating a fuzzy reasoning strategy and ANN based on the Adaptive Network to handle the demand containing inadequate knowledge. During testing, they obtained MAPE: 18% on average for some products. For unpredictable customer demands, neural methods have provided a robust forecasting strategy in a multi-level SC framework. A Greedy Aggregation Decomposition (GAD) approach is a generic approach to self-development in a discontinuous time-series forecasting method that considers double-based causes of variation, addressing a practical discontinuous issue of forecasting demand [81]. With a limited dataset, they outperformed SBA, Croston’s method, TSB, MA-7, SES, MAPA, MA-3, ADIDA, iADIDA, N-7 with a MAPE accuracy rate of 5.9%. [150] offered the SHEnSVM (Selective & Heterogeneous Ensemble of SVMs) model for sales forecasts. Individual SVMs were trained using samples produced by the bootstrap method, and grid search created parameters, as stated by the specified model. The optimum specific combo strategy was found using a genetic algorithm. They claimed a 10% increment using the SVM algorithm & a 64% on average enhancement in MAPE. The authors used beer data from three product variants in their tests. [74] integrated Deep-Learning (DL) technique, the SVR algorithm, and best-performing time-series analytic models utilizing the boosting ensemble approach to demand forecasting systems. Their DL implementation in the new integration strategy (MAPE: 24.7%) lowers mean forecasting error in the SC, outperforming both the conventional best-performing forecasting model (MAPE: 42.4%) and the unique integration strategy without DL (MAPE: 25.8%). XGBoost, ARIMA, and Snaive STL decomposition have outperformed solo and hybrid models and the modeling mix and have provided the best forecasting accuracy [53].
@@
10 most popular ML regressors, the EGD-SNet framework includes 11 dimensionality reduction techniques and 13 alternative feature selection algorithms. By locating the best hyperparameters, it employs Particle Swarm Optimizer (PSO) to train regressors intelligently. Also, it can create an end-to-end pipeline by selecting the right regressor, feature, and dimensionality reduction methodologies to accurately anticipate energy generation or demand for a specific geographical data set, depending on the features of the data. [8] implemented many DL methods, including data collecting, de-noising or pre-processing, feature extraction, and classification stages. Two primary DL models determine feature extraction. The first variation used three RNN structures: LSTM, BiLSTM, and GRU. The second variant used the temporal convolutional network (TCN). They used SoftMax, RT, RF, KNN, ANN, and SVM classifiers for an online dataset. TCN predicts COVID-19-restricted shipping risk almost 100% accurately.
@@
As the daily fish demand forecasting models for grocery merchants to reduce food waste and enhance sustainable SCs, [92] investigated LSTM, Feedforward NNs, Support Vector Regression, RF, and a Holt-Winters statistical model. The findings showed that the LSTM model provided the best outcomes in terms of root mean squared error (27.82), mean absolute error (20.63), and mean positive error (17.86). [54] forecasted solar Global Horizontal Irradiance using statistical and Deep Learning architectures, which aids grid management and power distribution and highlights Pakistan’s solar power potential in addressing global climate change. They employed SARIMAX, Prophet, LSTM, convolutional NN (CNN), and ANN statistical approaches. Error measures like $R^{2}$, MAE, MSE, and RMSE were used to evaluate each model’s performance. They concluded that SARIMAX and Prophet are ideal for long-term forecasts, whereas ANN, CNN, and LSTM are best for short-term forecasts. [118] found that the optimum model for every participant in the supply chain across all three inventory replenishment strategies is a stacked ensemble model consisting of XG Boost, Ada Boost, and Random Forest. According to a methodology for comparing forecasting methods developed by [120], the MLP method has a little edge over the CNN, LSTM, and CNN-LSTM approaches. [135] utilized data from the last five years to estimate demand for eight dairy products from five dairy production facilities using a direct multistep prediction method. ARIMA works effectively on a narrow subset of unpredictable series, whereas LSTMs excel at anticipating seasonal patterns. It outperforms ARIMA for trends. Monthly data decreased model training error.
@@
For the purpose of forecasting daily energy use, [108] investigated the effectiveness of three machine learning models, SVR, RF, and XGBoost; three deep learning models, RNNs, LSTM, and GRU; and ARIMA. For both very short-term load forecasting (VSTLF) and short-term load forecasting (STLF), the suggested XGBoost models beat competing models; the ARIMA model did the worst. [10] presented a smart platform for data-driven blood bank management that forecasts blood demand and balances blood collection and distribution based on optimal blood inventory management to avoid blood wastage and shortage. This improves blood quality and quantity, increasing blood collection by 11% and reducing blood waste by 20%. Balancing blood collection and distribution based on good blood inventory management and arranging blood donation sessions to avoid cancellations may lower inventory levels. [99] proposed a CNN-LSTM model with Swish Activation to estimate a store’s supply based on prior sale. This outperforms Rectified Linear Unit, the most effective activation function (ReLU). They forecasted sales using Multilayer Perceptron, LSTM cells, and CNNs. CNN-LSTM Model has a reduced RMSE, according to the experiment. Pharmaceutical businesses can use Shallow NN and DNN demand forecasting models for eight anatomical treatment chemical thematic drug groups [105]. Shallow NN models performed well for five of eight medication categories, while the ARIMA model performed best for the other three.
@@
EGD–SNet Framework: 10 ML regressors
@@
LSTM, Feedforward NNs, Support Vector Regression, RF, Holt–Winters model
@@
Supply chain inventory replenishment
@@
CNN–LSTM model with Swish Activation
@@
E–commerce product demand forecasting
@@
Adaptive Neuro–Fuzzy Inference System (ANFIS)
@@
Real–time e–order arrivals forecasting
@@
RF, XGBoost, Gradient boosting, AdaBoost, ANN, Hybrid RF–XGBoost–LR
@@
Based on the mentioned studies, we suggest considering the following recently best-performing hybrid time-series demand forecasting ML models: 1. XGBoost-LSTM [142] 2. FB-Prophet [76] 3. XGBoost-LightGBM [147] 4. M-GAN-XGBoost [139] 5. SARIMA, integrated AttConvLSTM, and FB-Prophet [138] 6. AUG-NN [66].
@@ 5.1 Information flow
Forecasting decisions affect further SC planning. As such, information flows across multiple phases of the SC process. Logistics superiority and better stock level synchronization are possible through a flow of demand information from downstream members to the upstream ones and the flow of production plan and delivery information from the upstream members to the downstream ones [132]. Like the three types of forecasting, there are three types of decisions in SC, strategic, operational, and tactical [124]. Recent studies have shown how the findings from one level can affect other decisions and limit the number of options for subsequent decisions [131].
Through forecasting, it is possible to attain efficiency by properly allocating resources in different areas such as workforce, capacity, inventory management, etc. There are need-based, supply-based, and demand-based models for forecasting and planning in such areas of SCM [114]. Each method requires some sort of information flow. The forecasted amount can be used to estimate the number of dependent inventory demands. The final product’s demand indirectly affects the required workforce, capacity, warehouse planning, and lower costs through optimization.

### 07 Feb 2024 v5

@@ 4.4 Exploratory data analysis (EDA) & data reduction (DR)
In this process, the target or dependent column and independent features are obtained. The DR, EDA, and clustering techniques reduce runtime and space during the deep-learning modeling phase. DR can be employed to decrease the size of a dataset while still keeping the data’s original integrity. In our framework, we suggest performing Feature Selection and Feature Extraction simultaneously after selecting the target column and finding redundant features; then, Discretisation may be performed if necessary. Then, the dataset will be ready for further analysis and model training.
@@ 4.4.1 Identifying redundant features
Covariance and correlation: In statistics, covariance refers to the amount that two features or factors change in tandem whose value lies in the (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) range. Positive covariance indicates they move in the same direction. Negative covariance means that any features are greater than the mean, and others are less than the mean, and vice-versa. Zero covariance means features may be independent under a certain hypothesis. On the other hand, correlation analysis is a widely used dimensionless measurement ranging from $-1$ to $+1$ to discover redundancies in numerical features that evaluate and quantify the relationship intensity. The features are positively correlated for correlation values greater than zero (0); for zero, they are independent; for less than zero, they are negatively correlated [46]. Covariance and correlation are directly proportional to each other. In numeric feature selection, correlation is better to use, as correlation analysis is scaled [-1, 1], but the covariance range is indefinite (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). We should choose correlation for better interpretation. Changes in location, size, or scale have no effect on correlation. However, both of them are limited to only being able to identify linearity.
$\chi^{2}$ correlation: The $\chi^{2}$ (Chi-Square) test is often used when dealing with nominal features and finite value sets. We can use the $\chi^{2}$ test to see whether there is any link between the values of two nominal features, where a probability table with joint events is established. If $\alpha$ (significance level) is less than the estimated one (or the $\chi^{2}$-value (calculated) ¿ table value), the null hypothesis gets discarded, and the two features can be said to be correlated statistically [46]. SC Analysts must remember that the $\chi^{2}$-test does not tell much about the strength of the relationship between two features. The $\chi^{2}$-test offers advantages such as resilience regarding data distribution, computational simplicity, extensive information produced from the test, utilization for investigations where parametric criteria cannot be satisfied, and scalability in processing data from two and multiple-group research. The drawbacks are sample size constraints and difficulty in comprehension when there are many (¿20) features.
@@ 4.4.2 Feature selection (FS)
  • In certain circumstances, FS results will still include significant important features that may obstruct the use of complicated training strategies.
Leading methods: In order to create FS techniques by combining a feature evaluation score and a cutting criterion, [3] recommended that functions based on information principle produce better accuracy, not suggesting any universal cutting condition. However, those independent of the metric perform best, and outcomes differ across models. For each kind of model, wrapper techniques were recommended to avoid this effect.
@@
We can use supervised FS if the data has class labels; otherwise, unsupervised FS is the best option. This approach generally maximizes clustering efficiency or the FS based on correlation, feature dependency, and priority. The primary premise is to eliminate features that bring almost no value beyond what is already provided by the existing features in the system. [94] suggested using feature dependency/similarity to reduce redundancy without needing a search procedure. An information compression metric called the maximum information compression index governs the clustering partitioning process, which uses features as the measure of similarity. Forward orthogonal search (FOS) is another unsupervised FS approach that aims to maximize the total reliance on the data to find relevant features [141]. Without compromising performance in clustering, the unsupervised FS used the Random Cluster Ensemble framework to compress the set of features by roughly 1/100 of its initial dimensions [32]. When compared to well-known classifications, precision/recall analyses revealed that feature weighting was highly successful in discovering the most suitable clusters [96].
@@ 4.4.3 Feature extraction
Principal component analysis (PCA): It aims to analyze a collection of features’ variance-covariance patterns employing a few linear combinations and seeks the optimal $k$ number of $n-dimensional$ orthogonal vectors for data description, where $k\geq n$. Accounting for the most critical percentage of the discrepancy in the original dataset, the principal component (the first derived feature) is produced in decreasing order of contribution. Typically, for containing $\geq$ 95% variance, just the top few principal components are retained. PCA is beneficial when many independent variables correlate with one another [70]. The principal component is quick and comprehensive and ensures a solution is found for all datasets [44].
@@ 4.5.2 Model fit and train
Two factors determine the proportions of these three sets: the number of data samples and training models. Some models require significant training data; therefore, the model should be tuned for more extensive training sets in this scenario. Models with fewer hyperparameters will be easier to validate and tune, allowing a small validation set size. An extensive validation set will benefit if the model contains more significant hyperparameters. There will be no requirement for a validation set if the model has no hyper-parameters or is challenging to adjust.
When using k-fold CV, the train-test dataset splitting is repeated for k-times, with each new set being given a shot at becoming the hold-out set. Time-series data cannot be used with k-fold CV directly since they believe there is no connection between the rows and that they are all separate instances. For time-series data, instances’ time horizon prevents arbitrarily dividing them into clusters. Instead, data should be segmented, and the chronological sequence of instances maintained. The term backtesting is used in time-series forecasting to describe the technique of evaluating models using past data. In meteorology, this is regarded as ’hindcasting’ rather than ’forecasting.’
@@ 4.5.3 Hyperparameter tuning
  • It enhances the opportunity to reproduce the ML process
@@ 4.5.5 Top forecasting models
Adaptive Neuro-Fuzzy Inference System (ANFIS) [55]
@@
For the demand forecasting procedure, [149] evaluated statistical models, RBFNN (Radial Basis Function NNs), and winter models with SVM. According to their conclusion, the efficiency of SVM outperforms other algorithms by about a mean MAPE outcomes threshold of 7.7%. [30] demonstrated a new AI-utilized forecasting approach evaluating a fuzzy reasoning strategy and ANN based on the Adaptive Network to handle the demand containing inadequate knowledge. During testing, they obtained MAPE 18% on average for some products. For unpredictable customer demands, neural methods have provided a robust forecasting strategy in a multi-level SC framework. A Greedy Aggregation Decomposition (GAD) approach is a generic approach to self-development in a discontinuous time-series forecasting method that considers double-based causes of variation, addressing a practical discontinuous issue of forecasting demand [81]. With a limited dataset, they outperformed SBA, Croston’s method, TSB, MA-7, SES, MAPA, MA-3, ADIDA, iADIDA, and N-7 with a MAPE accuracy rate of 5.9%. [150] offered the SHEnSVM (Selective & Heterogeneous Ensemble of SVMs) model for sales forecasts. Individual SVMs were trained using samples produced by the bootstrap method, and grid search created parameters, as stated by the specified model. The optimum specific combo strategy was found using a genetic algorithm. They claimed a 10% increment using the SVM algorithm & a 64% on average enhancement in MAPE. The authors used beer data from three product variants in their tests. [74] integrated DL technique, the SVR algorithm, and best-performing time-series analytic models utilizing the boosting ensemble approach to demand forecasting systems. Their DL implementation in the new integration strategy (MAPE: 24.7%) lowers mean forecasting error in the SC, outperforming both the conventional best-performing forecasting model (MAPE: 42.4%) and the unique integration strategy without DL (MAPE: 25.8%). XGBoost, ARIMA, and Snaive STL decomposition have outperformed solo and hybrid models and the modeling mix and have provided the best forecasting accuracy [53].
@@
The 10 most popular ML regressors in the EGD-SNet framework include 11 dimensionality reduction techniques and 13 alternative feature selection algorithms. It employs a Particle Swarm Optimizer (PSO) to train regressors intelligently by locating the best hyperparameters. Also, it can create an end-to-end pipeline by selecting the right regressor, feature, and dimensionality reduction methodologies to accurately anticipate energy generation or demand for a specific geographical data set, depending on the features of the data. [8] implemented many DL methods, including data collecting, de-noising or pre-processing, feature extraction, and classification stages. Two primary DL models determine feature extraction. The first variation used three RNN structures: LSTM, BiLSTM, and GRU. The second variant used the temporal convolutional network (TCN). They used SoftMax, RT, RF, KNN, ANN, and SVM classifiers for an online dataset. TCN predicts COVID-19-restricted shipping risk almost 100% accurately.
@@
As the daily fish demand forecasting models for grocery merchants to reduce food waste and enhance sustainable SCs, [92] investigated LSTM, Feedforward NNs, Support Vector Regression, RF, and a Holt-Winters statistical model. The findings showed that the LSTM model provided the best outcomes in terms of root mean squared error (27.82), mean absolute error (20.63), and mean positive error (17.86). [54] forecasted solar Global Horizontal Irradiance using statistical and Deep Learning architectures, which aids grid management and power distribution and highlights Pakistan’s solar power potential in addressing global climate change. They employed SARIMAX, Prophet, LSTM, convolutional NN (CNN), and ANN statistical approaches. Error measures like $R^{2}$, MAE, MSE, and RMSE were used to evaluate each model’s performance. They concluded that SARIMAX and Prophet are ideal for long-term forecasts, whereas ANN, CNN, and LSTM are best for short-term forecasts. [118] found that the optimum model for every participant in the SC across all three inventory replenishment strategies is a stacked ensemble model consisting of XG Boost, Ada Boost, and Random Forest. According to a methodology for comparing forecasting methods developed by [120], the MLP method has a little edge over the CNN, LSTM, and CNN-LSTM approaches. [135] utilized data from the last five years to estimate demand for eight dairy products from five dairy production facilities using a direct multistep prediction method. ARIMA works effectively on a narrow subset of unpredictable series, whereas LSTMs excel at anticipating seasonal patterns. It outperforms ARIMA for trends. Monthly data decreased model training error.
@@
For the purpose of forecasting daily energy use, [108] investigated the effectiveness of three ML models, SVR, RF, and XGBoost; three deep learning models, RNNs, LSTM, and GRU; and ARIMA. For both very short-term load forecasting (VSTLF) and short-term load forecasting (STLF), the suggested XGBoost models beat competing models; the ARIMA model did the worst. [10] presented a smart platform for data-driven blood bank management that forecasts blood demand and balances blood collection and distribution based on optimal blood inventory management to avoid blood wastage and shortage. This improves blood quality and quantity, increasing blood collection by 11% and reducing blood waste by 20%. Balancing blood collection and distribution based on good blood inventory management and arranging blood donation sessions to avoid cancellations may lower inventory levels. [99] proposed a CNN-LSTM model with Swish Activation to estimate a store’s supply based on prior sale. This outperforms Rectified Linear Unit (ReLU), the most effective activation function. They forecasted sales using Multilayer Perceptron, LSTM cells, and CNNs. CNN-LSTM Model has a reduced RMSE, according to the experiment. Pharmaceutical businesses can use Shallow NN and DNN demand forecasting models for eight anatomical treatment chemical thematic drug groups [105]. Shallow NN models performed well for five of eight medication categories, while the ARIMA model performed best for the other three.
@@
EGD-SNet Framework: 10 ML regressors
@@
LSTM, Feedforward NNs, Support Vector Regression, RF, Holt-Winters model
@@
SC inventory replenishment
@@
CNN-LSTM model with Swish Activation
@@
E-commerce product demand forecasting
@@
Adaptive Neuro-Fuzzy Inference System (ANFIS)
@@
Real-time e-order arrivals forecasting
@@
RF, XGBoost, Gradient boosting, AdaBoost, ANN, Hybrid RF-XGBoost-LR
@@
Based on the mentioned studies, we suggest considering the following recently best-performing hybrid time-series demand forecasting ML models: 1. XGBoost-LSTM [142] 2. FB-Prophet [76] 3. XGBoost-LightGBM [147] 4. M-GAN-XGBoost [139] 5. SARIMA integrated AttConvLSTM, and FB-Prophet [138] 6. AUG-NN [66].
@@ 5.1 Information flow
Forecasting decisions affect further SC planning. As such, information flows across multiple phases of the SC process. Logistics superiority and better stock level synchronization are possible through a flow of demand information from downstream members to the upstream ones and the flow of production plan and delivery information from the upstream members to the downstream ones [132]. Like the three types of forecasting, there are three types of decisions in SC: strategic, operational, and tactical [124]. Recent studies have shown how the findings from one level can affect other decisions and limit the number of options for subsequent decisions [131].
It is possible to attain efficiency through forecasting by properly allocating resources in different areas such as workforce, capacity, inventory management, etc. There are need-based, supply-based, and demand-based models for forecasting and planning in such areas of SCM [114]. Each method requires some sort of information flow. The forecasted amount can be used to estimate the number of dependent inventory demands. The final product’s demand indirectly affects the required workforce, capacity, warehouse planning, and lower costs through optimization.